through:(4,2), parallel to y= - 3/4x-5
Answer by mananth(16949)

y = - 3/4 x -5
Divide by 1
y = - 3/4 x -5
Compare this equation with y=mx+b
slope m = - 3/4
The slope of a line parallel to the above line will be the same
The slope of the required line will be - 3/4
m= - 3/4 ,point ( 4 , 2 )
Find b by plugging the values of m & the point in
y=mx+b
2 = -3 + b
b= 5
m= - 3/4
Plug value of the slope and b in y = mx +b
The required equation is y = - 3/4 x + 5

m.ananth@hotmail.ca